OpenStudy (anonymous):

Solve -x^2+8x-7 and explain how? :)

OpenStudy (anonymous):

is it =0

OpenStudy (anonymous):

Multiply the equation by -1 to make x positive

OpenStudy (anonymous):

Yes it is =0 sorry!

OpenStudy (anonymous):

Multiply the equation by -1

OpenStudy (anonymous):

X^2-8x+7=0 What multiplies to 7 and ads to -8?

OpenStudy (anonymous):

We get x^2-8x+7=0

OpenStudy (anonymous):

Okay so then it'd be x^2-8x+7=0, then what?

OpenStudy (anonymous):

@oliviaellen23 i m working it out

OpenStudy (anonymous):

(X-7)(x-1)

OpenStudy (anonymous):

Oh so you just factor! Okay! Thank you guys! :)

OpenStudy (anonymous):

x^2-7x-1x+7=0 x(x-7)-1(x-7)=0 (x-1)(x-7)=0 So either (x-1)=0 ------> Therefore x = +1 or (x-7)=0 -------------> Therefore x= +7

OpenStudy (anonymous):

It Says solve so we have to find the value of x not just factor the equation

OpenStudy (anonymous):

(X-7)=0. (X-1)=0 Solve for x

OpenStudy (anonymous):

If u don't understand anything let me know.